OpenStudy (anonymous):

I need to solve y=2^x using logarithmic differentiation.

OpenStudy (shamil98):

\[(a^x)' = a^x(\ln~a)\]

OpenStudy (shamil98):

Would it not be just: \[y' = 2^x(\ln 2)\] ?

OpenStudy (anonymous):

I got the answer using that method but need this method. I would take log of both sides but I get stuck

OpenStudy (anonymous):

heres where I lock up